Circular No. 7459

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMETS C/1999 G5, C/1999 H7, C/1999 J12, C/1999 X2, C/2000 N2

Further to IAUC 7458, D. Hammer provides measurements of four comets found by M. Oates in archival images obtained in 1999 with the SOHO C2 telescope only and of one comet found by M. Meyer in real-time SOHO C3 telescope data. The discovery observations for these Kreutz sungrazers are given below, while MPEC 2000-O10 contains the fully reduced observations and orbital elements (by B. G. Marsden).

1999 UT  α_{2000}  δ_{2000}  Comet
Apr. 11.246  1^{h}24.^{m}4 + 7^{o}16′ C/1999 G5
20.413  1 57.3 +10 21  C/1999 H7
May 3.035  2 43.2 +14 10  C/1999 J12
Dec. 3.896 16 42.2 −23 41  C/1999 X2
2000 UT  α_{2000}  δ_{2000}  Comet
July 8.221  7^{h}00.^{m}0 +20^{o}51′ C/2000 N2
2000 EB_{173}

D. Rabinowitz, Physics Department, Yale University, reports, on behalf of the QUEST collaboration (cf. Schaefer et al. 1999, Ap.J. 524, L103), the following photometry of the plutino 2000 EB_{173} (cf. MPEC 2000-L09, MPO 1583): Mar. 1–15 UT, V = 19.86 ± 0.06, R = 19.27 ± 0.03 (average of the discovery and prediscovery observations on seven nights with the 1-m Schmidt telescope, Llano del Hato, Venezuela); June 3, R = 19.40 ± 0.12 (Yale 1.0-m telescope, Cerro Tololo); June 10, B = 21.10 ± 0.11, V = 20.11 ± 0.08, R = 19.51 ± 0.06, I = 19.13 ± 0.07 (2.4-m MDM telescope, Kitt Peak); June 15, V = 20.18 ± 0.03, R = 19.63 ± 0.03 (3.5-m WIYN telescope, Kitt Peak). He adds: “A summed 45-min exposure in V with the WIYN telescope reveals a stellar intensity profile in 1′′.3 seeing and no brightness variations > 3 percent over 1.25 hr. The March observations show no variability greater than ∼ 10 percent. M. Brown (California Institute of Technology) notes a flat and featureless reflectance spectrum from 1.4 to 2.4 μm from 10-m Keck observations on June 18 and 19.”

SUPERNOVA 2000cs IN MCG +7-34-15 Corrigendum. On IAUC 7446, line 8, for Osoyoss, read Osoyoos,

2000 July 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green